a car starts from rest and moves with uniform acceleration of alpha m/s^2 along a straight line . it then retards uniformly at a rate of beta m/s^2 and stops. if t is the time elapsed find the average speed

Grade:10

1 Answers

Arun
6 years ago
Dear bhavan
 
Let maximum velocity be x 
then x/α+x/β=t 
Solving, 
x=t/(1/α+1/β) 
=αβt/(α+β)
 
Regards
Arun (askIITians forum expert)